# The Best Instruction-Tuning Data are Those That Fit

### v2

@@ 5 GRAPE-Picking From Real-World SFT Datasets
In this section, we leverage the findings from the earlier experiments and demonstrate the effectiveness of GRAPE to customize training data for each base model by selecting from available datasets with overlapping instructions. Here, we do not generate any new responses for the instructions; it only selects from existing ones. We evaluate GRAPE on the fully open dataset used in post-training phases of Tulu-3 (Lambert et al., 2024) and Olmo-2 (OLMo et al., 2025). The details are presented below.
@@ 5.1 Data Mixture Details
Tulu-3 (Lambert et al., 2024) is a fully open-source collection of post-training recipes, including supervised fine-tuning and preference alignment data. Olmo-2 (OLMo et al., 2025) is a fully open-source language model. Both Tulu-3 and Olmo-2 use the same data mixture during the supervised fine-tuning stage, but different data mixtures and source models for generating preference data for different sizes of their models: Tulu-3-8B/70B and Olmo-2-7B/13B. To demonstrate the effectiveness of GRAPE, we collected the overlapping instructions from both models and gather their corresponding responses. From the preference data, we retained only the winning responses. We formed our candidate pool with those instructions with at least two distinct responses, resulting in a dataset of 350.4k unique instructions and about 1.03 million total instruction-response pairs for evaluation with GRAPE. We do not apply further processing of these data or any filtering on top of GRAPE.
@@ 5.2 Evaluation
We evaluate on a set of commonly used benchmarks spanning over coding, math, knowledge and instruction-following. We evaluated on LeetCode (Guo et al., 2024a), MATH (Hendrycks et al., 2021b), BigBenchHard(BBH) (Suzgun et al., 2022), MMLU (Hendrycks et al., 2021a), and AlpacaEval-V2 (Dubois et al., 2024). LeetCode, MATH, BBH and MMLU are evaluated the in the same way as in (Yuan et al., 2024), where we use zero-shot for MATH and MMLU, 3-shot example for BBH. We use the same AlpacaEval-v2 as in OpenInstruct.
@@ 5.3 Baselines
  • Responses From The Original SFT on The Subset. In this experiment, we keep the same set of instructions (350.4k) and pair each instruction with the response from the SFT mixture.
  • Random Response Candidate From The Pool. Randomly pairing each instruction with a candidate response for that instruction, which may come from either the SFT mixture or a winning response in the preference-learning data.
  • Lowest-Probability Response. Instead of taking the highest probability response to each instruction for the base model, we do the opposite on and take the lowest for each.
  • Entire SFT Dataset. We train on the entire SFT data mixture of Tulu with 939k instances.
  • All Available Responses For The Overlapping Instructions. In this setting, we train on the entire candidate pool for all overlapping instructions, resulting in 103.6k instances.
  • S2L (Yang et al., 2024b). S2L is a state-of-the-art unsupervised data selection baseline designed to select balanced subsets from large datasets by leveraging training loss trajectories. It first trains a small reference model within the same model family and record loss trajectory for each training example, then applies K-means clustering on the loss trajectory and samples equally from each. We use Llama-3.2-1B as reference for Llama3.1-8B and Qwen2.5-0.5B for Qwen2.5-7B. For Mistral-v0.3-7B, we use itself as reference since no smaller models in the family are available. For S2L, we select the same number of data from the same pool as GRAPE. Further details in Appendix B.
  • All data available for SFT mixture. We train with all the available responses from all these datasets that share prompts with the SFT mixture. Which gives approximately 1.58 million (4.5 times thesize of GRAPE-curated dataset). This is the union of the entire dataset and GRAPE’s candidate pool.
@@
Table 2: GRAPE on the Tulu-Olmo collection. For Llama3.1-8B base model, we included Tulu3-SFT model’s results. The “*”-marked number for MATH is obtained using 4-shot prompting. We train all the models (except that we took Tulu3-SFT numbers directly) for 1 epoch with a learning rate of $10^{-5}$.
@@ 5.4 Results
As shown in Table 2, models fine-tuned on responses selected by GRAPE outperforms the strong baselines we constructed, especially the one that trains over all available data by significant margins across the 3 models. Remarkably, Using roughly 1/6 training computation (Tulu3-8B-SFT was trained for 2 epochs on 3 times of data), our performance exceeds that of Tulu3-8B-SFT. Also, GRAPE outperforms state-of-the-art data-selection approaches like S2L, despite its simplicity and efficiency, further highlighting its effectiveness in diverse real-world scenarios and strengthening its overall practicability for large-scale instruction tuning and robust generalization. Without the need to synthesize any new data, one can easily leverage established datasets sourced from the web to customize a dataset for each base model that yields better fine-tuning outcome.
These results feature GRAPE not only as an effective strategy to enhance performances, but a handy approach to improve fine-tuning efficiency.
5.5 Further Experiments
Item
Metric Data

### v3

@@ 5 GRAPE-Picking From Real-World SFT Datasets
In this section, we leverage the findings from the earlier experiments and demonstrate the effectiveness of GRAPE to customize training data for each base model by selecting from available datasets with overlapping instructions. Here, we do not generate any new responses for the instructions; it only selects from existing ones. We evaluate GRAPE on the fully open dataset used in post-training phases of Tulu-3 (Lambert et al., 2024) and Olmo-2 (OLMo et al., 2025). The details are presented below. We discuss additional results in Appendix B
@@ 5.1 Data Mixture Details
Tulu-3 (Lambert et al., 2024) is a fully open-source collection of post-training recipes, including supervised fine-tuning and preference alignment data. Olmo-2 (OLMo et al., 2025) is a fully open-source language model. Both Tulu-3 and Olmo-2 use the same data mixture during the supervised fine-tuning stage, but different data mixtures and source models for generating preference data for different sizes of their models: Tulu-3-8B/70B and Olmo-2-7B/13B. To demonstrate the effectiveness of GRAPE, we collected the overlapping instructions from both models and gather their corresponding responses.
Figure 4: Breakdown of GRAPE-selected responses for 1K Tulu instructions vary significantly across base models, reflecting its highly model-oriented nature over responses. Details in Appendix H
From the preference data, we retained only the winning responses. We formed our candidate pool with those instructions with at least two distinct responses, resulting in a dataset of 350.4K unique instructions and about 1.03 million total instruction-response pairs for evaluation with GRAPE. We do not apply further processing of these data or any filtering on top of GRAPE.
@@ 5.2 Evaluation
We evaluate on a set of commonly used benchmarks spanning over coding, math, knowledge and instruction-following. We evaluated on LeetCode (Guo et al., 2024a), MATH (Hendrycks et al., 2021b), BigBenchHard(BBH) (Suzgun et al., 2022), MMLU (Hendrycks et al., 2021a), and AlpacaEval-V2 (Dubois et al., 2024). LeetCode, MATH, BBH and MMLU are evaluated the in the same way as in (Yuan et al., 2024b), where we use zero-shot for MATH and MMLU, 3-shot example for BBH. We use the same AlpacaEval-v2 as in OpenInstruct.
@@ 5.3 Baselines
We extensively compare against three baseline types: controlled baselines with fixed instructions, scaling baselines with increased data, and state-of-the-art selection methods. Additional details in Appendix C
Controlled Baselines
We include three baselines to isolate the effect of GRAPE’s response selection. SFT-only replaces GRAPE-selected responses with those from the original SFT dataset, pairing each instruction with a standard reference response to measure improvement over presumably good SFT responses. Random selects candidate responses uniformly at random from the pool for the same set of instructions, establishing a noise-tolerant baseline. Reverse-GRAPE instead selects responses with the highest perplexity. We test if responses diverging the most from the base model’s distribution degrade performance, providing a contrast that sharpens the effectiveness of GRAPE.
Scaling Baselines
[…]
State-Of-The-Art SFT Data Selection Approaches
We compare GRAPE against recent state-of-the-art data selection methods. LESS (Xia et al., 2024) selects data based on influence scores on validation tasks. We follow the implementation setup from Dai et al. (Dai et al., 2025). Emb-NV selects training data close to validation in embedding space using NV-Embed-V2 (Lee et al., 2025), following (Ivison et al., 2025). S2L (Yang et al., 2024b) clusters data via loss trajectories from small reference models (Llama-3.2-1B, Qwen2.5-0.5B, Mistral-v0.3-7B) and samples uniformly across clusters to match GRAPE’s data budget.
@@
[…]
@@ 5.4 Results
GRAPE
Qwen2.5
-72B
Llama3.1
-405B
Gemma
-it-9B
LC 28.1 25.8 16.3 26.9
WR 33.1 24.3 17.0 20.2
Table 3: Alpaca-Eval2 On Magpie-Zoo Xu et al. (2024c).
As shown in Table 2, models fine-tuned on responses selected by GRAPE outperforms the strong baselines we constructed, especially the one that trains over all available data by significant margins across the 3 models.
Remarkably, using roughly 1/6 training computation (Tulu3-8B-SFT was trained for 2 epochs on 3 times of data), our performance exceeds that of Tulu3-8B-SFT.
@@
Also, GRAPE outperforms state-of-the-art data-selection approaches like S2L, despite its simplicity and efficiency, further highlighting its effectiveness in diverse real-world scenarios and making it a practical option for real-world SFT setups with minimal data engineering effort. Without the need to synthesize any new data, one can easily leverage established datasets sourced from the web to customize a dataset for each base model that yields better fine-tuning outcome.
These results highlight GRAPE as an effective and efficient selection strategy for real-world SFT.
5.5 Ablations
Remains Effective Even with a Single Response Source
In Section 4, we demonstrated how GRAPE enables practitioners to refine model-generated responses for improved training outcomes. Beyond that, GRAPE can optimize responses from a single generator. The Magpie-Zoo (Xu et al., 2024c) dataset contains a fixed set of instructions and multiple versions of response sets each generated by different language models. Using the Magpie-Zoo instruction set, we sample 10 responses per instruction from Qwen2.5-72B-Instruct, select in-distribution responses with GRAPE, to train a Mistral-v0.3-7B model. We compare with replicas that achieve top-3 performance from Magpie-Zoo.
Model Data Avg.
Self-Distilled 28.4(-)
Mistral-7B Original-UI 32.7
Self-Distilled 29.4(-)
Llama3.1-8B Original-UI 36.8
Self-Distilled 15.1(-)
Llama3.2-3B Original-UI 20.3
Table 4: Performance Degradation From Self-Distillation On UI.
As shown in Table 3, GRAPE-selected responses further boost the performance. Given that batch sampling from a strong model introduces minimal latency (Zhong et al., 2024; Zhou et al., 2024e), this result positions GRAPE as a practical and efficient data curation strategy—achieving strong results with just a single generator and no additional engineering overhead.
Why GRAPE Outperforms Self-Generated Responses
Dataset Model Response Generator N Acc.
MATH Mistral Llama3.1-70B 10 18.2
MATH Mistral FT-Mistral 10 15.9 (-)
MATH Llemma Llama3.1-70B 10 26.2
MATH Llemma FT-Llemma 10 23.6 (-)
MATH Mistral $M$-AnsAug 22.3
MATH Mistral FT-Mistral 10 20.6 (-)
MATH Llemma $M$-AnsAug 28.1
MATH Llemma FT-Llemma 10 21.4 (-)
Table 5: Self-generation on MATH dataset. FT-Mistral refers to the model right above that row finetuned from either $M$-AnsAug or Llama3.1-70B-Instruct produced solutions. N stands for the number of responses sampled.
To probe the effectiveness of GRAPE, we ablate it against a degenerate alternative: self-generation, where the model is fine-tuned on its own outputs.
We fine-tune a base model using responses generated by its previously fine-tuned variant. This setup consistently degrades performance (Table 4). We further confirm this on the MATH dataset (Hendrycks et al., 2021b): responses from strong models like Llama3.1-70B-Instruct or MetaMathQA (Yu et al., 2024) are used to fine-tune a model that then generates new solutions for another round of fine-tuning. Again, performance drops (Table 5).
This failure arises from distributional collapse (see § 2.2): self-generated responses become increasingly narrow and repetitive, reinforcing biases and reducing exposure to diverse reasoning. Correctness alone is insufficient—external diversity is essential for generalization.
GRAPE avoids this collapse by selecting external responses that are both diverse and distribution-aligned, preserving semantic breadth and stylistic variability while staying true to the model’s pretraining. This enables more stable and generalizable fine-tuning.
6 Conclusion
We present GRAPE, a simple yet effective method for improving supervised fine-tuning by selecting responses aligned with the base model’s pretrained distribution. GRAPE requires only a forward pass over candidate responses, making it highly efficient and easy to integrate. Despite its simplicity, GRAPE consistently outperforms stronger baselines using significantly larger datasets and surpasses more complex, costly data selection methods. Our study affirms that carefully aligning SFT data with a model’s pretrained distribution yields substantial performance and efficiency gains.
7 Discussion and Limitations
Response versus Instance Level Selection
GRAPE selects responses: it begins with a fixed set of instructions and evaluates multiple candidate responses for each. This differs from instance-level selection approaches (e.g., Kung et al. (2023); Li et al. (2024c); Wang et al. (2024b)), which focus on choosing which instructions to include based on factors such as coverage, skill-balancing or difficulty. In contrast, GRAPE focuses on the quality of supervision—that is, selecting responses to provide the most effective learning signal. Importantly, this response-centric perspective is complementary, not contradictory, to instance-level or instruction-based selection methods; both might be combined to enhance overall data quality and training efficiency.
Limitations
Like many other data selection algorithms (Du et al., 2023; Li et al., 2024c; Xia et al., 2024; Das & Khetan, 2023; Kang et al., 2024; Mekala et al., 2024; Yang et al., 2024c; Zhang et al., 2024b; Pan et al., 2024; Dai et al., 2025; Ivison et al., 2025; Bhatt et al., 2024; Yin & Rush, 2024; Liu et al., 2024b), GRAPE assumes a quality-controlled candidate pool from which to select samples. Furthermore, because GRAPE relies on the base model itself, its selection effectiveness may be influenced by the model’s inherent capabilities.
8 Acknowledgment
This project is partly supported by NSF under award No. 2019897. This research used the DeltaAI advanced computing and data resource, which is supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. DeltaAI is a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See experiments in Section 4 and 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See section 5.6, where we show pursuing in-distribution answers in the wrong way can lead to performance degradations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate ”Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See appendices.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use publicly available datasets and models. Our method only requires computing the normalized probability of training data, which can be easily done with any open-sourced machine learning codebase.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See Experiments Sections and Appendices.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We follow standard evaluation paradigms by computing pass@1 accuracy with greedy sampling, which does not explicitly involve randomness.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer ”Yes” if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: See appendicies, where we provide information about GPU resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: See paper.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper presents work whose goal is to advance the field of Machine Learning. It investigates fundamental aspects of instruction-tuning of language models and should not have direct societal impacts or implications that should be discussed here specifically, to the best of the authors’ knowledge.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not release data or models that have a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Yes, we cite all the datasets and pretrained models used in the paper, which are all open-sourced for research use.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not have crowdsourcing experiments or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not have crowdsourcing experiments or research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not involve LLMs as a core component in implementing or developing the method of this work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Parameter Distance
We measure the L2 norm of model parameter difference between fine-tuned and pre-trained checkpoints, as a signal of how much the distribution has drifted during SFT ichi Amari (2016); Cover & Thomas (2006). We notice that training over well-matched distribution shifts the parameter less than training over those ill-matched.
Mistral-7B-v0.3 Llama3.1-8B Qwen2.5
GRAPE 8.006 8.196 8.426
Worst 8.029 8.202 8.467
Table 6: Performance comparison across different models
Appendix B Further Experiments
B.1 Comparing with Reward Based Selection
We compare GRAPE with purely reward-based selection. where for each instruction, we select the response with the highest scalar reward as determined by a reward model - Skywork-Reward-Llama3.1-8B-v0.2. Once the top-ranked response for each instruction is selected, we proceed with standard supervised fine-tuning on the resulting instruction-response pairs.The RFT setup provides a natural contrast to our proposed GRAPE method by emphasizing reward alignment over base-model alignment, thereby enabling us to disentangle the effects of distribution matching versus reward optimization in SFT data selection.
As shown in Table 7, GRAPE outperforms reward-based selection across both models and all benchmarks. These results suggest that aligning supervision with the base model’s own distributional preferences—rather than relying on external reward models—can yield better task performance.
Model Method AE WR AE WR (LC) LeetCode MATH MMLU BBH Avg
LLaMA3.1-8B GRAPE 15.2 14.8 19.4 32.1 64.5 69.6 35.9
Reward 14.0 14.5 17.2 31.3 63.3 69.0 34.9
Mistral-v0.3-7B GRAPE 13.9 13.6 18.3 24.2 59.2 62.3 31.9
Reward 12.5 13.9 13.3 22.4 58.5 62.2 30.5
Table 7: Performance comparison between GRAPE and reward-based selection across benchmarks for LLaMA3.1-8B and Mistral-v0.3-7B. Metrics are benchmark-specific scores (higher is better).
B.2 Experiment on OpenHermes
Item Metric Data